Three points with position vectors, b and c are said to be colinear. If the parallelogram with adjacent sides a - b and a - c has zero geometry area. Use this fact to check whether or not the following triples of points are collinear


(a) (2,2,3), (6,1,5) (2,4,3)


(b) (2,3,3), (3,7,5), (0,-5,-1)


(c) (1,3,2), (4,2,1), (1,0,2)


1
Expert's answer
2022-05-20T08:14:39-0400

(a)


"\vec{a}-\vec{b}=(-4, 1, -2)"

"\vec{a}-\vec{c}=(0, -2, 0)"

"(\vec{a}-\vec{b})\times (\vec{a}-\vec{c})=\begin{vmatrix}
 \vec{i} & \vec{j} & \vec{k} \\
 -4 & 1 & -2 \\
0 & -2 & 0
\end{vmatrix}"

"=2\begin{vmatrix}
 \vec{i} & \vec{k} \\
 -4 & -2
\end{vmatrix}=-4\vec{i}+8\vec{k}"

"\big|(\vec{a}-\vec{b})\times (\vec{a}-\vec{c})\big|=\sqrt{(-4)^2+(8)^2}=4\sqrt{5}\not=0"

Three points are not collinear.


(b)


"\vec{a}-\vec{b}=(-1, -4, -2)"

"\vec{a}-\vec{c}=(2, 8, 4)"

"(\vec{a}-\vec{b})\times (\vec{a}-\vec{c})=\begin{vmatrix}
 \vec{i} & \vec{j} & \vec{k} \\
 -1 & -4 & -2 \\
2 & 8 & 4
\end{vmatrix}"

"=2\begin{vmatrix}
 \vec{j} & \vec{k} \\
 -4 & -2
\end{vmatrix}-8\begin{vmatrix}
 \vec{i} & \vec{k} \\
 -1 & -2
\end{vmatrix}+4\begin{vmatrix}
 \vec{i} & \vec{j} \\
 -1 & -4
\end{vmatrix}"


"=-4\vec{j}+8\vec{k}+16\vec{i}-8\vec{k}-16\vec{i}+4\vec{j}=\vec{0}"


"\big|(\vec{a}-\vec{b})\times (\vec{a}-\vec{c})\big|=\sqrt{(-4)^2+(8)^2}=0"

Three points are collinear.


(c)


"\vec{a}-\vec{b}=(-3, 1, 1)"

"\vec{a}-\vec{c}=(0, -3, 0)"

"(\vec{a}-\vec{b})\times (\vec{a}-\vec{c})=\begin{vmatrix}
 \vec{i} & \vec{j} & \vec{k} \\
 -3 & 1 &1 \\
0 & -3 & 0
\end{vmatrix}"

"=-3\begin{vmatrix}
 \vec{i} & \vec{k} \\
 -3 & 1
\end{vmatrix}=-3\vec{i}-9\vec{k}"




"\big|(\vec{a}-\vec{b})\times (\vec{a}-\vec{c})\big|=\sqrt{(-3)^2+(-9)^2}"

"=3\sqrt{10}\not=0"

Three points are not collinear.